OpenStudy (anonymous):

Solve by using the perfect squares method. x2 + 16x + 64= 0 A. 8 B. 16 C. –8 D. –16

OpenStudy (anonymous):

C -8

OpenStudy (anonymous):

(x+8)^2=0 sqrt{(x+8)^2}=sqrt{0} x+8=0 x=-8

OpenStudy (anonymous):

Any guess what would be first step ??

OpenStudy (zale101):

use the quadratic formula and plug in the numbers that replaces ax^2+bx+c=0 as x^2 + 16x + 64= 0 so a=1 b=16 and c=64 quadratic formula \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] now plug and chug!